# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ 4 Simulation Studies
We conducted a simulation study to evaluate the performance of these methods under the context of survey observational data in estimating various population-level estimands, particularly under different degrees of overlap and model misspecification. Below we detail our simulation design and key findings.
4.1 Simulating design
@@
Our data-generating process was adapted from the multi-stage strategy employed by Austin et al. and Lenis et al.. [11, 13] We simulated a population of $1,000,000$ individuals, organized into 10 strata $(j=1,\ldots,10)$. Each stratum was further divided into 20 clusters, resulting in a total of 200 clusters $(k=1,\ldots,200)$, with 5000 individuals in each cluster $(i=1,\ldots,5000)$. For each individual, we generated six baseline covariates $(l=1,\ldots,6)$ from normal distributions, where the vector of covariates for the $i$-th individual is denoted as $\mathbf{X}_{i}=\{X_{1i},X_{2i},X_{3i},X_{4i},X_{5i},X_{6i}\}$. The mean of these covariates were allowed to vary at both the stratum and cluster levels. For the $l$-th covariate, a stratum-specific random effect for the $j$ th stratum $\nu_{l,j}^{\text{stratum }}\sim N\left(0,\sigma_{l}^{\text{stratum }}\right)$ was generated, along with a cluster-specific random effect for the $k$th cluster $\nu_{l,k}^{\text{cluster }}\sim N\left(0,\sigma_{l}^{\text{cluster }}\right)$. The $l$-th covariate for the $i$-th individual in the $j$-th stratum and $k$th cluster was then generated as $X_{l,ijk}\sim N\left(\nu_{l,j}^{\text{stratum }}+\nu_{l,k}^{\text{cluster }},1\right)$. We set the standard deviation associated with stratum-level means to $\sigma^{\text{stratum }}=0.35$ and with cluster-level means to $\sigma^{\text{cluster }}=0.15$.
We then generated a binary treatment assignment for each individual using a logistic model to determine the probability of treatment:
$\operatorname{logit}\left(e_{\text{sp}}(\mathbf{X}_{i})\right)=a_{0}+\psi\left(\boldsymbol{a}_{1:6}^{T}\mathbf{X}_{i}+a_{7}X_{1i}X_{2i}\right)$
@@
where $\boldsymbol{a}_{1:6}=(\log(1.1),\log(1.25),\log(1.5),\log(1.75),\log(2),\log(2.5))^{T}$, and $a_{7}=\log(1.1)$. Treatment status $Z_{i}$ was simulated from a Bernoulli distribution: $Z_{i}\sim\operatorname{Bernoulli}(e_{\text{sp}}(\mathbf{X}_{i}))$. The scalar $\psi$ reflects the strength of confounding and dictates the level of overlap in the propensity score. [35] We set $\psi=0.6$ or $2$ to mimic scenarios with good or poor overlap in the full population, as visualized in Fig. 1(a) and Fig. 1(b), respectively. For both scenarios, the overall treatment prevalence was maintained at approximately 30% by adjusting $a_{0}$ to $\log(35/80)$ for the good overlap and $\log(20/80)$ for the poor overlap, allowing us to isolate the impact of overlap when evaluating estimator performance.
(a) good overlap scenario
(b) poor overlap scenario
Figure 1: Distribution of Propensity Scores in Different Overlap Scenarios
@@
We simulated continuous outcomes for each individual $i$ in the population using a linear combination of covariates, treatment effects, and interaction terms. Specifically, the potential outcome $Y_{i}$ was modeled as:
$Y_{i}=b_{0}+\delta_{0}\left(\mathbf{b}^{T}\mathbf{X}_{i}+b_{7}X_{1i}X_{2i}\right)+Z_{i}\left(\delta_{1}+\delta_{2}\left(\mathbf{b}^{T}\mathbf{X}_{i}+b_{8}X_{1i}X_{2i}\right)\right)+\varepsilon$
@@
where $\varepsilon\sim N(0,1)$. We set $b_{0}=0$ and $\boldsymbol{b}=(2.5,-2,1.75,-1.25,1.5,1.1)^{T}$ for the covariate coefficients. The average treatment effect was represented by $\delta_{1}=1$, while $\delta_{0}=0.3$ and $\delta_{2}=0.2$ introduced heterogeneity in the control and treatment conditions, respectively. Additionally, two interaction terms were included, with $b_{7}=2.5$ and $b_{8}=1.5$, to capture the more complex influence of baseline covariates on outcomes across both conditions.
[…]
@@
To validate the robustness of our estimator under different sampling mechanisms, we further extended the simulation by introducing an alternative sampling mechanism dependent on both baseline covariates and treatment assignment. We used logistic regression to model the sampling probability for individual $i$ as a function of both the treatment assignment indicator $Z_{i}$ and covariates $\mathbf{X}_{i}$, allowing sampling to reflect more complex real-world dependencies. The sampling probability $p_{z,i}=P(S_{i}=1\mid Z_{i},\mathbf{X}_{i})$ was computed as $\operatorname{logit}(p_{z,i})=c_{0}+\delta_{z}^{(s)}Z_{i}+\mathbf{c}^{T}\mathbf{X}_{i}$, where $c_{0}=\log\left(\frac{0.005}{1-0.005}\right)$ defines the baseline sampling rate of 0.5% consistent with the above multi-stage design, and $\delta_{z}^{(s)}=\log(0.9)$ slightly reduces the sampling probability for the treatment group ($Z=1$), addressing possible treatment-related biases. The vector $\mathbf{c}=(c_{1},c_{2},c_{3},c_{4},c_{5},c_{6})^{T}$ represents the covariate coefficients with values $c_{1}=\log(1.05),c_{2}=\log(1.10),c_{3}=\log(1.15),c_{4}=\log(1.10),c_{5}=\log(1.05),$ and $c_{6}=\log(1.10)$, corresponding to the influences of covariates $X_{1}$ through $X_{6}$ on the selection probability. To avoid the impact of near-zero values leading to extreme weights in further estimation, we will perform and recommend truncation with $\alpha=0.001$ on the fitted propensity scores after incorporating the varied survey weights in this settings. The sampling indicator $S_{i}$ was subsequently generated as a Bernoulli random variable with probability $p_{z,i}$ for individual $i$, capturing the interaction between treatment assignment and covariates in determining sample inclusion. More details and results of this alternative simulation are presented in Web Appendix B.2.
@@
For each sample within each scenario, we evaluated four types of estimators: the PSW (propensity score weighting estimator) and three augmentation methods, referred to as MOM (moment estimator), CVR (clever covariate regression estimator), and WET (Weighted Regression estimator), as previously described. To investigate the role of survey weights in estimating the propensity scores, we primarily compare two strategies for integrating survey weights in the propensity score model: as model weights (W.PS) and as an additional covariate (C.PS). The latter involves directly adjusting for survey weight as one additional single covariate in the propensity score model, which is distinct from including the balancing weight as a clever covariate in the outcome regression model. Additionally, results for unweighted propensity scores (U.PS) and survey weights applied as both model weights and covariates (CW.PS) are presented in the Web Appendix B.1. These comparisons may shed light on previous debates in the literature regarding which level of covariate balance should we achieve with propensity score and in what way.[3, 4]
@@
Given the complexity of the design, it is instructive to note the total number of considered scenarios. We examined two levels of overlap (good vs. poor), two primary methods for incorporating survey weights into the propensity score model (W.PS vs. C.PS), four estimators (PSW, MOM, CVR, WET), three estimands (PATE, PATT, PATO), and four model specification scenarios (Cor|Cor, Mis|Cor, Cor|Mis, Mis|Mis). This combination yields $2\times 2\times 4\times 3\times 4=192$ distinct configurations in the main analysis. Additional scenarios that incorporate the unweighted and combined weighting-plus-covariate propensity score approaches and the sampling mechanism depending on treatment increase this number even further. The factorial nature of this design ensures that our conclusions are drawn from a broad range of realistic conditions, and the resulting comparisons are adequately comprehensive. Table 4 outlines these factors and their levels, providing an overview of the multiplicity of scenarios explored.
@@
Performance metrics included relative bias, relative efficiency, and empirical coverage of the 95% confidence intervals. Relative bias was calculated as the percentage difference between the estimated and true population treatment effects, averaged across 1,000 simulations. Relative efficiency was determined as the ratio of the Monte Carlo variance of the reference estimator (the balancing weight estimator for PATE with W.PS under correct model specification) to that of each evaluated estimator. Empirical coverage was assessed by calculating the proportion of confidence intervals, constructed using sandwich variances, that contained the true population treatment effect. The simulation code and implementations of the survey-weighted balancing weight estimator and the augmented estimators for PATE, PATT, and PATO are available online at https://github.com/ykzeng-yale/SW_PSW, and they are implemented in the PSweight R package (Version 2.1.0).
@@ 4.2 Simulations results
In this section, we focus on the primary four types of estimators (PSW, MOM, CVR and WET) and two distinct methods of incorporating survey weights into propensity score models (W.PS and C.PS). Table 5 presents the simulation results under the good overlap scenario, summarizing the relative bias, relative efficiency, and coverage of PATE, PATT, and PATO under various modeling strategies with either correctly or incorrectly specified propensity score and outcome regression models. In parallel, Table 6 presents the results when the overlap is poor. In addition to these primary results, we present more comprehensive analyses that include U.PS and CW.PS in Web Tables 3 to 5 and 6 to 8 for each estimand under good and poor overlap. Furthermore, simulation results for the extended scenario, where sampling depends on treatment assignment, are provided in Web Tables 9 and 10 for W.PS and C.PS results, and Web Tables 11 to 13 and 14 to 16 for more comprehensive analyses. To facilitate the summary of findings, we focus on the following three thorny questions; that is, how survey weights should be incorporated at each stage of the balancing weight estimator, which augmented estimators perform better, and which method is more robust (i.e., under poor overlap and model misspecification).
@@ 4.2.1 How should survey weights be incorporated into the propensity score weighting estimator?
Focusing first on the results from the weighting estimator alone (PSW), our simulation results support our theoretical conclusion that incorporating survey weights in both the propensity score and outcome estimation leads to unbiased and efficient propensity score weighting estimators of the population-level treatment effect. Under good overlap (Table 5), using survey-weighted propensity score models (W.PS) consistently produces lower bias, higher relative efficiency, and coverage closer to the nominal 95% compared to using survey weights as an additional covariate (C.PS). This pattern is particularly evident for PATO, where W.PS attains near-ideal balance and performance across all specification scenarios. These observations, shown in the “Cor|Cor” columns of the first two rows of each estimand corresponding to PSW results in Table 5, indicate that W.PS excels when the propensity score model is correct and continues to outperform C.PS even when misspecified, as seen in the ”Mis|Cor” columns.
@@
Under poor overlap (Table 6), the advantage of W.PS over C.PS becomes even more pronounced. While C.PS suffers from inflated bias and poor coverage, particularly for PATE and PATT, W.PS maintains lower bias, higher efficiency, and closer-to-nominal coverage across all estimands. Particularly, inference for PATO is more robust to poor overlap, PSW with W.PS achieves the lowest biases and highest relative efficiency among all PSW methods, underscoring its ability to attain population-level covariate balance. Moreover, PSW with W.PS for PATO maintains near-nominal coverage at 93.8%, while C.PS drops below 90%. Even when the propensity score model is misspecified and interaction terms are ignored (“Mis|Cor” columns in Tables 5 and 6), W.PS consistently outperforms C.PS in terms of bias and coverage, demonstrating its desirability and robustness under model misspecification, regardless of overlap.
@@
Additional evidence from alternative simulation results (Web Tables 9 and 10), where sampling depends on treatment assignment, confirms W.PS’s advantages. In these extended scenarios, W.PS achieves near-zero bias, while C.PS’s bias deteriorates to 25%–54% under good overlap (Web Table 9). Extreme survey weights, likely when sampling depends on treatment, exacerbate model extrapolation and misspecification issues under C.PS. In contrast, W.PS handles these conditions effectively, maintaining robust performance. Inference for PATO demonstrates outstanding robustness in the alternative sampling mechanism, especially under poor overlap (Web Table 10). PSW with W.PS for PATO achieves the smallest bias (0.5%) and the highest efficiency—nearly 30 times the reference method (PSW with W.PS for PATE) —and is the only PSW estimator maintaining nominal coverage at 94.4%. These findings align with insights from Ridgeway et al.,[4] which highlight the benefits of survey-weighted propensity scores across various sampling mechanisms, particularly when influenced by treatment assignment, though limited to PATT estimation. Our results extend this by demonstrating W.PS’s robustness across various estimands (PATE, PATT, and PATO) and under different overlap and misspecification conditions.
@@
While initially demonstrated with simple weighting estimators (PSW), these advantages generalize to augmented estimators. In our main scenarios with multi-stage sampling setups (Tables 5 and 6), augmented estimators with W.PS or C.PS perform similarly; this indicates that augmentation provides some protective effect and confers a degree of robustness regardless of how the propensity scores are estimated. However, under treatment-dependent sampling (Web Tables 9 and 10), the differences become substantial, and W.PS wins over C.PS in both bias and coverage even with augmentation. Thus, our simulations highlight W.PS as the preferred method for incorporating survey weights into the propensity score model, regardless of whether an outcome regression component is used in the final estimation.
@@
4.2.2 Is there a universal winner among the augmented estimators when all models are correctly specified?

### v2

@@ 4 Simulation Studies
4.1 Simulation design
@@
Our data-generating process was adapted from the multi-stage strategy employed by Austin et al. and Lenis et al. [11, 13] We simulated a population of $1,000,000$ individuals, organized into 10 strata $(j=1,\ldots,10)$. Each stratum was further divided into 20 clusters, resulting in a total of 200 clusters indexed globally as $(k=1,\ldots,200)$, with 5000 individuals in each cluster $(i=1,\ldots,5000)$. For each individual, we generated six baseline covariates $(l=1,\ldots,6)$ from normal distributions, where the vector of covariates for the $i$-th individual is denoted as $\mathbf{X}_{i}=\{X_{1i},X_{2i},X_{3i},X_{4i},X_{5i},X_{6i}\}$. The means of these covariates were allowed to vary at both the stratum and cluster levels. For the $l$-th covariate, a stratum-specific random effect for the $j$ th stratum $\nu_{l,j}^{\text{stratum }}\sim N\left(0,\sigma_{l}^{\text{stratum }}\right)$ was generated, along with a cluster-specific random effect for the $k$ -th cluster $\nu_{l,k}^{\text{cluster }}\sim N\left(0,\sigma_{l}^{\text{cluster }}\right)$. The $l$-th covariate for the $i$-th individual in the $j$-th stratum and $k$ -th cluster was then generated as $X_{l,ijk}\sim N\left(\nu_{l,j}^{\text{stratum }}+\nu_{l,k}^{\text{cluster }},1\right)$. This hierarchical structure builds in correlation within clusters (through shared stratum and cluster effects) as well as within strata (through shared stratum effects only), while maintaining independence between strata. We set $\sigma^{\text{stratum }}=0.35$ and $\sigma^{\text{cluster }}=0.15$ throughout.
We specified the true population-level propensity score for each individual using a logistic model:
$\operatorname{logit}\left(e_{\text{sp}}(\mathbf{X}_{i})\right)=a_{0}+\psi\left(\boldsymbol{a}_{1:6}^{T}\mathbf{X}_{i}+a_{7}X_{1i}X_{2i}\right),$
@@
where $\boldsymbol{a}_{1:6}=(\log(1.1),\log(1.25),\log(1.5),\log(1.75),\log(2),\log(2.5))^{T}$, and $a_{7}=\log(1.1)$. Treatment assignment $Z_{i}$ was simulated from a Bernoulli distribution: $Z_{i}\sim\operatorname{Bernoulli}(e_{\text{sp}}(\mathbf{X}_{i}))$. The scalar $\psi$ reflects the strength of confounding and dictates the level of overlap in the propensity score. We set $\psi=0.6$ or $2$ to mimic scenarios with good or poor overlap in the full population, as visualized in Web Figure 1(a) and Web Figure 1(b), respectively. For both scenarios, the overall treatment prevalence was maintained at approximately 30% by adjusting $a_{0}$ to $\log(35/80)$ for the good overlap and $\log(20/80)$ for the poor overlap, allowing us to isolate the impact of overlap when evaluating estimator performance.
@@
We simulated continuous outcomes for each individual $i$ in the population using a linear combination of covariates, treatment effects, and interaction terms, with
$Y_{i}=b_{0}+\delta_{0}\left(\mathbf{b}^{T}\mathbf{X}_{i}+b_{7}X_{1i}X_{2i}\right)+Z_{i}\left(\delta_{1}+\delta_{2}\left(\mathbf{b}^{T}\mathbf{X}_{i}+b_{8}X_{1i}X_{2i}\right)\right)+\varepsilon,$
@@
where $\varepsilon\sim N(0,1)$. We set $b_{0}=0$ and $\boldsymbol{b}=(2.5,-2,1.75,-1.25,1.5,1.1)^{T}$. The average treatment effect was represented by $\delta_{1}=1$, while $\delta_{0}=0.3$ and $\delta_{2}=0.2$ introduced heterogeneity in the control and treatment conditions, respectively. Additionally, two interaction terms were included, with $b_{7}=2.5$ and $b_{8}=1.5$, to capture the more complex influence of baseline covariates on outcomes across both conditions.
We drew 1,000 random samples with 5000 individuals each (representing 0.5% of the overall population) first under a prospective framework using a multi-stage sampling design consistent with Austin et al.[11] and Lenis et al.,[13], where the sampling mechanism was independent of the treatment assignment mechanism. Sample sizes were allocated across the 10 strata as $\{850,750,700,650,600,400,350,300,250,150\}$, ensuring that the observed sample would have a different distribution from the population. Within each stratum, five clusters were selected using simple random sampling, and from each selected cluster, an equal number of individuals were sampled. For each sampled individual, we calculated the true survey weights by taking the inverse of the sampling probability, reflecting the number of individuals in the population that the sampled individual represents.
@@
To further assess the performance of our estimators under retrospective designs, we extended the simulation by introducing an alternative sampling mechanism which now depends on both baseline covariates and treatment assignment. We used logistic regression to model the sampling probability for individual $i$ as a function of both the treatment assignment indicator $Z_{i}$ and covariates $\mathbf{X}_{i}$, allowing sampling to reflect more complex real-world dependencies. The sampling probability $p_{z,i}=P(S_{i}=1\mid Z_{i},\mathbf{X}_{i})$ was computed as $\operatorname{logit}(p_{z,i})=c_{0}+\delta_{z}^{(s)}Z_{i}+\mathbf{c}^{T}\mathbf{X}_{i}$, where $c_{0}=\log\left(\frac{0.005}{1-0.005}\right)$ defines the baseline sampling rate of 0.5% consistent with the above multi-stage design, and $\delta_{z}^{(s)}=\log(0.9)$ slightly reduces the sampling probability for the treatment group. The coefficient vector was specified as $\mathbf{c}=(c_{1},c_{2},c_{3},c_{4},c_{5},c_{6})^{T}=(\log(1.05),\log(1.10),\log(1.15),\log(1.10),\log(1.05),\log(1.10))^{T}$. The sampling indicator $S_{i}$ was subsequently generated as a Bernoulli random variable with probability $p_{z,i}$ for individual $i$, capturing the interaction between treatment assignment and covariates in determining sample inclusion. More details and results of this alternative simulation are presented in Web Appendix B.2.
@@
For each sample within each scenario, we evaluated four types of estimators: the PSW (propensity score weighting estimator) and three augmentation methods, referred to as MOM, CVR and WET. To investigate the role of survey weights in estimating the propensity scores, we primarily compare two strategies for integrating survey weights in the propensity score model: as model weights (W.PS) and as an additional covariate (C.PS). The latter involves directly adjusting for survey weight as one additional single covariate in the propensity score model, which is distinct from including the balancing weight as a clever covariate in the outcome regression model. Additionally, results for unweighted propensity scores (U.PS) and survey weights applied as both model weights and covariates (CW.PS) are presented in the Web Appendix B.1. These comparisons may shed light on previous debates in the literature regarding which level of covariate balance should we achieve with propensity score and in what way.[3, 4] For estimating the sample-level propensity score $e_{\text{fp}}(\mathbf{X})$ used in approximating $p(\mathbf{X}_{i})$ for the augmented estimators, we use standard unweighted logistic regression as an effective approximation under this rare sampling condition for the theoretically correct nonlinear logistic form in Equation (1)). See Web Appendix B.3.1 for additional sensitivity analyses.
@@
Given the complexity of the design, it is instructive to note the total number of considered scenarios. We examined two levels of overlap (good vs. poor), two primary methods for incorporating survey weights into the propensity score model (W.PS vs. C.PS), four estimators (PSW, MOM, CVR, WET), three estimands (PATE, PATT, PATO), and four model specification scenarios (Cor|Cor, Mis|Cor, Cor|Mis, Mis|Mis). This combination yields $2\times 2\times 4\times 3\times 4=192$ distinct configurations in the main analysis. Additional scenarios that incorporate the unweighted and combined weighting-plus-covariate propensity score approaches and the sampling mechanism depending on treatment increase this number even further. The factorial nature of this design ensures that our conclusions are drawn from a broad range of realistic conditions, and the resulting comparisons are adequately comprehensive. Table 4 outlines these factors and their levels for a quick reference.
@@
Beyond the main simulation framework, we conducted sensitivity analyses to assess the robustness of our proposed estimators under alternative configurations. In Web Appendix B.3, we systematically evaluate two aspects: (1) the choice between the principled approach using standard unweighted logistic regression versus maximum likelihood estimation for the theoretically correct nonlinear logistic form of the sample-level propensity score $e_{\text{fp}}(\mathbf{X})$ under retrospective treatment-dependent sampling (Section B.3.1), and (2) robustness across diverse sampling mechanisms including varying treatment-dependent sampling strengths (odds ratios of $p_{1}(\mathbf{x})/p_{0}(\mathbf{x})$ from 0.5 to 2.0), complex $Z$-$X$ interactions in the true sampling model, and alternative link functions such as probit (Section B.3.2). These sensitivity analyses demonstrate the robust performance of our proposed estimators under possible model misspecification and varying complex sampling designs.
Performance metrics included relative bias, relative efficiency, and empirical coverage of the 95% confidence intervals. Relative bias was calculated as the percentage difference between the estimated and true population treatment effects, averaged across 1,000 simulations. Relative efficiency was determined as the ratio of the Monte Carlo variance of the reference estimator (the balancing weight estimator for PATE with W.PS under correct model specification) to that of each evaluated estimator. Empirical coverage was assessed by calculating the proportion of confidence intervals, constructed using sandwich variance estimators, that contained the true population treatment effect. The simulation code and implementations of the survey-weighted balancing weight estimator and the augmented estimators for PATE, PATT, and PATO are available online at https://github.com/ykzeng-yale/SW_PSW, and they are implemented in the PSweight R package (Version 2.1.0).
@@ 4.2 Simulations results
In this section, we focus on the primary four types of estimators (PSW, MOM, CVR and WET) and two distinct methods of incorporating survey weights into propensity score models (W.PS and C.PS). Table 5 presents the simulation results under the good overlap scenario, summarizing the relative bias, relative efficiency, and coverage of PATE, PATT, and PATO under various modeling strategies with either correctly or incorrectly specified propensity score and outcome regression models. In parallel, Table 6 presents the results when the overlap is poor. In addition to these primary results, we present more comprehensive analyses that include U.PS and CW.PS in Web Tables 3 to 5 and 6 to 8 for each estimand under good and poor overlap scenarios. Furthermore, simulation results for the extended scenario, where sampling depends on treatment assignment, are provided in Web Tables 9 and 10 for W.PS and C.PS results, and Web Tables 11 to 13 and 14 to 16 for more comprehensive analyses. To facilitate the summary of findings, we focus on the following three thorny questions; that is, how survey weights should be incorporated at each stage of the balancing weight estimator, which augmented estimators perform better, and which method is empirically more robust (i.e., under poor overlap and model misspecification).
@@ 4.2.1 How should survey weights be incorporated into the propensity score weighting estimator?
Focusing first on the results from the weighting estimator alone (PSW), our simulation results support our theoretical conclusion that incorporating survey weights in both the propensity score estimation and final standardization step leads to unbiased estimation of the population-level treatment effect. Under good overlap (Table 5), using survey-weighted propensity score models (W.PS) consistently produces lower bias, higher relative efficiency, and coverage closer to the nominal 95% compared to using survey weights as an additional covariate (C.PS). This pattern is particularly evident for PATO, where W.PS shows satisfactory performance across all specification scenarios. Overall, the results indicate that W.PS excels when the propensity score model is correct and continues to outperform C.PS even under a misspecified propensity score model (the ”Mis|Cor” columns).
@@
Under poor overlap (Table 6), the advantage of W.PS over C.PS becomes more pronounced. While C.PS suffers from inflated bias and poor coverage, particularly for estimating PATE and PATT, W.PS maintains lower bias, higher efficiency, and closer-to-nominal coverage across all estimands. It is worth noting that inference for PATO is more robust to poor overlap, as PSW with W.PS achieves the lowest biases and highest relative efficiency among all PSW methods. Moreover, PSW with W.PS for PATO maintains near-nominal coverage at 93.8%, while C.PS drops below 90%. Even when the propensity score model is misspecified and interaction terms are ignored (“Mis|Cor” columns in Tables 5 and 6), W.PS consistently outperforms C.PS in terms of bias and coverage.
@@
Additional empirical evidence from additional simulations (Web Tables 9 and 10), where sampling depends on treatment assignment, lead to the same findings. In these extended scenarios, W.PS achieves near-zero bias, while the relative bias of C.PS deteriorates to 25%–54% even under good overlap (Web Table 9). Extreme survey weights, likely when sampling depends on treatment, exacerbate model extrapolation and misspecification issues for C.PS. Inference for PATO demonstrates outstanding robustness in the alternative sampling mechanism, especially under poor overlap (Web Table 10). Importantly, PSW with W.PS for PATO achieves the smallest bias (0.5%) and the highest efficiency—nearly 30 times the reference method (PSW with W.PS for PATE) —and is the only PSW estimator maintaining nominal coverage (94.4%) under this challenging scenario. These findings align with insights from Ridgeway et al.,[4] which highlight the benefits of survey-weighted propensity scores across various sampling mechanisms, though their previous investigation was restricted to PATT. Our results cover additional estimands (PATE, PATT, and PATO) different overlap and misspecification scenarios.
@@
The above findings also apply to augmented estimators with some nuances. In our main simulation scenarios (Tables 5 and 6), augmented estimators with W.PS or C.PS perform similarly; this indicates that augmentation provides some protective effect and confers a degree of robustness regardless of how the propensity scores are estimated. However, under treatment-dependent sampling (Web Tables 9 and 10), the differences become substantial, and W.PS wins over C.PS in both bias and coverage even with augmentation. Overall, our empirical results highlight W.PS as the preferred method for incorporating survey weights into propensity score estimation, regardless of whether an outcome regression component is used to construct the final causal effect estimator.
@@
4.2.2 Is there a global winner among the augmented estimators when all models are correctly specified?